Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report page - 20130208

Observers: Nancy Elias-Rosa, Antonia Morales-Garoffolo
Data reduction team: Sandra Benitez, Stefan Taubenberger

Log

Observations with EFOSC2. Humidity between 50-70%. The seeing has been oscillating all the night.

Calibration data done in the afternoon.

UT 23:51      skyflats  UBVRi

UT 00:22      Focus + image analysis

--DIMM SEEING 1.0"--

UT 00:49      s82_0222+0054                  airmass=1.66     gri                  Photometric STD in gunn filters

UT 00:55      GD71                           airmass=1.42     gr11,13,16 sl 1.0    Flux STD

--DIMM SEEING 1.1"-- 

UT 01:17      LSQ12hxg                       airmass=1.19     gr13 sl 1.0

UT 03:06      OGLE-2012-SN-040               airmass=1.35     gr13 sl 1.0          Spectra aborted after first exposure when we realized it wasn't the SN. 

--DIMM SEEING 1.1" --

UT 04:12      LSQ13fn                        airmass=1.30     gr11 sl 1.0          

--DIMM SEEING 1.6"--

UT 06:07      L745a                          airmass=1.37     gr11,13,16 sl 1.5

--DIMM SEEING 1.8"--

UT 06:25     MASTEROTJ093953.18+165516.4     airmass=1.56     gr13 sl 1.5          **Classification: young Ia SN

--DIMM SEEING 2.2"!!--

UT 06:47     PSNJ152113475-0722183           airmass=1.88     gr11 sl 1.5          The seeing is becoming even higher (2.5"). We decided to observe a brighter target.

UT 07:15     SN2012hs                        airmass=1.17     gr11 sl 1.5 

--DIMM SEEING 2.0"--

UT 08:45     EG274                           airmass=1.28     gr11,13,16 sl 1.5

--DIMM SEEING 1.7" and HUMIDITY 85%!--

UT 09:03     CLOSING by high humidity and condensation!

UT 09:10     END Of the night.

Good night.
   
Comments